# StickerConv: Generating Multimodal Empathetic Responses from Scratch

### 20 Jan 2024 v1

@@ 5.3 Result Analysis
From Table 1, PEGS shows remarkable performance in several metrics, particularly in EX (Explorations), DIFF-EPITOME and diversity measuresm indicating a superior ability in generating diverse and exploratory responses. However, PEGS’s lower scores in IP, ER could be due to two reasons: (1) The general metrics only evaluate the text output and underestimate the actual empathy ability of the PEGS. (2) In multimodal empathy dialogue task, stickers play a significant role in conveying emotional reactions and explanations, while text handles more explorative tasks.
Models Rank Emp-txt Emp-sticker Con-txt Con-sticker
Vicuna 2.12 3.6768 4.3221
ChatGLM3 1.99 3.6908 4.3405
PEGS 1.88 3.7354 4.2534 4.2739 4.1155
Table 2: Results of LLM-Scorer.
Table 2 reveals PEGS’s strong performance in ranking, empathy text (Emp-txt), empathy sticker, and consistency of sticker responses (Con-sticker). However, its lower consistency in text (Con-txt) might be attributed to its emphasis on empathetic stickers, which could compromise some textual context consistency. This is in line with the additional information suggesting that the empathetic effectiveness of stickers might require some trade-offs in textual context consistency.
Models MM-Rel Inter-MM-Rel CLIP-Score
PEGS 69.41 59.06 21.03
Table 3: Results of multimodal evaluate.
Table 3 showcases the multimodal capabilities of PEGS. PEGS performs well in multimodal relevance (MM-Rel and Inter-MM-Rel) and consistency (CLIP-Score). This suggests that PEGS effectively integrates different modalities in its responses, maintaining relevance and consistency across them.
@@ 6 Conclusion
We explored the concept of multimodal empathetic response and created the StickerConv, the first dataset specifically designed for multimodal empathetic dialogue. We developed the Agent for StickerConv, a sophisticated LLM-based multi-agent system capable of simulating human-like interactions using stickers, thereby creating multimodal empathetic responses. Building upon the StickerConv, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal dialogue model. This model is adept at both perceiving and generating stickers, effectively enhancing the communication experience. Furthermore, we established a comprehensive set of evaluation metrics for multimodal empathy tasks based on LLM. We are confident this work will be a valuable asset in advancing research in the field of multimodal empathetic dialogue systems.
Appendix A Appendix
A.1 Dataset Statistics
Table 4 offers a statistical breakdown of the dataset, segmented into training, validation, and test splits. It provides a quantitative overview, detailing the number of instances in each subset, the unique stickers, turn counts, and token counts.
@@ A.1 Dataset Statistics
Figure 5 showcases the emotion distribution within user profiles in the Agent for StickerConv. The chart illustrates the percentage share of each emotion, offering insight into the prevalence of affective expressions within the dataset.
@@
Appendix B Case Study
We conduct a series of case studies to provide a more straightforward illustration of the performance of our model instruction-tuned by the StickerConv in multimodal empatheric conversations. These cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers.
As shown in Figure 7, in the positive emotion condition, PEGS will reward users with praise and support. As shown in Figure 8, two conversations in a negative scenario are displayed. PEGS as a conciliator can comfort users and guide their emotions towards positivity.
Appendix C Prompts
@@ C.1 Sticker Process
C.2 Sticker Agent Chat
Figure 10 shows the chat prompt, the summary encompasses the persona of the user or system, the current status, and the core traits generated based on memory. Figure 11, 12, 13 show the process of used sticker.
C.3 LLM-Based Scorer
@@
Figures 15 and 16 depict the templates for the consistency scorer. The selection of the appropriate template depends on the modality of the response generated by the model. Specifically, Figure 15 is utilized for scoring purely textual responses. In contrast, when the response is multimodal, the scoring is conducted using the template shown in Figure 16.
Similarly, Figures 17 and 18 present the templates for the empathy scorer. The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
C.4 Joint Learning
Figure 19 shows the instruction prompts for joint learning. We employed an instruction template similar to Vicuna_delta_v0. For every query or response that uses a sticker, we append an <IMG> identifier at the end.
Figure 7: Examples of conversations with positive emotions. Users can chat with multimodal content (text and stickers) and will receive positive multimodal responses.
Figure 8: Examples of conversations with negative emotions. Our model can empathize with users who are suffering from negative emotions, such as sadness (left) or anger (right) . Users will be comforted and guided to positive sentiments.

### 16 Feb 2024 v2

@@ 6.3 Result Analysis
Table 1 reports the results for text metrics. PEGS’s performance on the Dist-n is slightly lower than that of the Vicuna. This may be attributed to PEGS expanding Vicuna’s vocabulary to facilitate sticker perception and generation, impacting PEGS’s text diversity to a certain degree. In our experiment, the training objective of tool learning includes both text generation and tool calling. These two objectives may interfere with each other. As a result, the tool model’s text capabilities fall short of the text fine-tuning model’s. PEGS’s end-to-end structure integrates multimodal inputs and outputs, streamlines tasks with unified training objectives, and yields the best text results. PEGS-Gen utilizes fewer special tokens (32), thereby outperforming PEGS-Ret and PEGS-RAG in text metrics. These findings corroborate the efficacy of PEGS framework in generating text responses of high quality and accuracy.
Model Freq. MMr. f-MMr.
Vicuna-tool 0.141 0.725 0.602
ChatGLM3-tool 0.905 0.659 0.647
PEGS-Ret 0.850 0.674 0.653
PEGS-RAG 0.847 0.680 0.659
PEGS-Gen 0.811 0.672 0.647
Table 2: Results of multimodal metrics.
Table 2 displays the results for multimodal metrics, where Freq represents the relative frequency of stickers replies from each model. Although Vicuna-tool achieves a high MMr (72.48), this result is largely due to its lower Freq (0.141), which may not accurately represent its multimodal interaction capabilities. F-MMr is more robust than MMr because it accounts for Freq. PEGS excels in f-MMr, showcasing that its end-to-end structure, integrating text and stickers, achieves high consistency in multimodal reply generation.
Model EMP-txt EMP-mm CON. Rank
Vicuna-text 3.677 - 4.322 4.527
Vicuna-tool 3.822 3.799 4.247 4.093
ChatGLM3-text 3.691 - 4.341 5.037
ChatGLM3-tool 3.700 3.760 4.220 4.400
PEGS-Ret 3.873 4.040 4.380 4.030
PEGS-RAG 3.927 4.076 4.370 3.900
PEGS-Gen 3.768 4.353 4.404 1.917
Table 3: Results of LLM-based metrics.
Table 3 presents the evaluation results for the LLM. Relative to the baseline model, the two tool-learning models outperform the text model in ranking. This underscores the pivotal role of stickers in enhancing empathetic communication. According to the results from LLM-based metrics, PEGS can produce empathetic texts while ensuring high consistency, and utilize stickers to enhance emotional resonance. Table 2 and PEGS-Gen’s EMP-mm (4.353) indicates that the quality of the multimodal response directly correlates with the extent to which it enhances empathy.
Model StiGQ. Et. Es. Con. Flu. Inf. StiD.
Vicuna-tool 4.09 4.07 3.78 4.08 4.23 4.08 3.20
ChatGLM3-tool 4.32 4.06 3.24 4.11 4.58 3.99 3.10
PEGS-Ret - 4.11 4.17 4.22 4.36 4.09 3.40
PEGS-RAG 3.37 4.12 2.92 4.22 4.36 4.10 3.80
PEGS-Gen 4.53 4.29 4.19 4.37 4.47 4.26 3.60
Table 4: Results of the human evaluation.
Table 4 shows the result of human evaluation. PEGS notably outstrips the two baseline models across most evaluative indicators, save for Flu, thus corroborating the efficacy of the PEGS framework. The Flu of PEGS ($<$4.5) is lower compared to ChatGLM3-tool (4.58), likely because its high Inf ($>$4.09) results in a slight reduction in Flu. According to Inf and Et, the amount of information is a crucial factor influencing empathy. PEGS-RAG excels in sticker diversity (3.8) but records the lowest score in sticker generation quality (3.37). This observation suggests that although the RAG strategy enhances sticker variety, it might also reduce the quality of sticker generation in this experiment. Analysis of StiGQ, Es, and StiD reveals that multimodal responses can enhance empathy, and this enhancement is positively correlated with the responses’ quality. When considering all evaluation of human metrics, PEGS-Gen stands out remarkably, partly because of its minimal use of special tokens and partly due to its proficient sticker generation capability (owe to end-to-end structure). These outcomes highlight PEGS-Gen’s exceptional ability to produce high-quality, emotionally resonant, and diverse multimodal responses.
Further analysis is presented in Appendix D. We collect the case study in Appendix E.
@@ 7 Conclusion
We explored the concept of multimodal empathetic response and created the STICKERCONV, the first dataset specifically designed for multimodal empathetic dialogue. We developed the Agent for STICKERCONV, a sophisticated LLM-based multi-agent system capable of simulating human-like interactions using stickers, thereby creating multimodal empathetic responses. Building upon the STICKERCONV, we developed PErceive and Generate Stickers (PEGS), an advanced multimodal empathetic dialogue framework. This framework adeptly perceives and generates stickers, effectively enhancing the communication experience. Furthermore, we established a comprehensive set of evaluation metrics for multimodal empathy tasks based on LLM. We are confident this work will be a valuable asset in advancing research in the field of multimodal empathetic dialogue systems.
Limitations
Although our study has achieved certain advancements, it is not without its limitations. Firstly, STICKERCONV is exclusively in English, implying that our model and evaluations might not be directly translatable to other linguistic contexts. Secondly, although we investigate the retrieval-augmented generation (RAG) variant of the PEGS, hoping to enhance the quality and diversity of sticker generation, the RAG version has demonstrated instability in generating quality stickers in practice. Despite these issues, the pursuit of diversity and quality in sticker generation means that the explorations into the RAG version remain valuable. It suggests a promising research trajectory: how to elevate the stability and variety of the generated content by refining algorithms or amalgamating additional technological solutions. Finally, from our experiments, PEGS and two baseline tool-model exhibit abnormal frequencies in the usage of stickers. Aligning the model’s sticker usage frequency and timing with human-like interaction patterns is an important area for future exploration.
Ethics Statement
The STICKERCONV dataset, produced by the large language model (LLM), therefore sidesteps privacy concerns. Nonetheless, LLMs may inherently harbor biases, and both STICKERCONV and PEGS could inadvertently magnify these biases in their application. Despite our efforts to mitigate these biases by incorporating a diverse range of personality information into Agent4SC, it is important to acknowledge that biases intrinsic to LLMs cannot be completely eradicated. This inherent bias is an ongoing challenge that needs to be addressed continuously through iterative improvement and monitoring of the model. PEGS reveals its robust capability in creating multimodal empathetic replies, adept at grasping human emotions and generating pertinent multimodal content. Such abilities, if exploited, could potentially manipulate user emotions. To forestall these risks, we have implemented a mechanism for detecting NSFW content in the generation of multimodal content. It is imperative to underscore that, despite our model’s proficiency in aiding the understanding and response to human emotions, it is not intended to entirely supplant human emotional communication. Excessive dependency on AI for empathetic interactions might erode human empathic abilities and the cultivation of emotional intelligence.
Appendix A Dataset Statistics
In our dataset creation, it is important to note that vector databases numbered 1 to 80 are exclusively employed for the creation of training and cross-validation sets, whereas those numbered 81 to 100 are reserved for the construction of the test set. We configure the agent’s "select" action’s top-K parameter to 10.
Table 5 offers a statistical breakdown of the dataset, segmented into training, validation, and test splits. It provides a quantitative overview, detailing the number of instances in each subset, the unique stickers, turn counts, and token counts. Our dataset is comparable in size to PhotoChat Zang et al. (2021). To our knowledge, it is the first multimodal empathetic dialogue dataset, uniquely integrating stickers as non-textual elements to enhance empathetic communication more effectively.
@@ Appendix A Dataset Statistics
Figure 4 shows the emotional distribution of stickers used by the User and the System. It reveals a striking trend: users have a significant preference for stickers that convey negative emotion, in contrast to the system. The system predominantly utilizes stickers to express neutral and positive emotion. This comparison not only reflects the distinct emotional expression preferences between users and the system but also highlights the system’s active and supportive role in interactions. Figure 10 and 11 present two examples from STICKERCONV.
Figure 4: The chart of emotional distribution in the choice of stickers between the User and the System.
Figure 5 showcases the emotion distribution within user profiles in the Agent for STICKERCONV. The chart illustrates the percentage share of each emotion, offering insight into the prevalence of affective expressions within the dataset. Increasing personality distributions enhances the Agent4SC’s diversity, thereby helping to mitigate the inherent bias issue in LLMs. Figure 9 shows a profile example across various emotions.
@@
Appendix B Implementation Details of Agent4SC
B.1 Tool Module
Before advancing to the follwing process, we altered the background of the SER30K stickers from black to white. The detailed process of building the Tool module is as follows:
Filter
Leveraging the LLaVa-v1.5-13b (Liu et al., 2023a), we craft a nuanced Chain of Thought (CoT) Wei et al. (2023) procedure to meticulously scrutinize each SER30K sticker. This approach entails the model initially summarizing the sticker’s content, subsequently assessing its relevance to casual conversation, and finally determining suitability through iterative responses. Each sticker undergoes multiple evaluations to ensure accuracy.
Annotation
Filtered stickers undergo a second CoT analysis to answer 5 questions about their content, emotions, usage contexts, tone, and any comedic or satirical elements. This detailed annotation process provides deep insights, structured as Q&A pairs for further use.
Knowledge
Leveraging ChatGPT-3.5-Turbo, we extract critical information from the Q&A pairs, generating a comprehensive understanding of each sticker, including its description, emotion, and usage recommendation.
Split
We segment the filtered dataset into smaller vector databases, incorporating the original emotional labels alongside freshly acquired knowledge. Each database maintains an emotional sticker distribution that aligns with that of the original dataset.
Following the process described, we obtain a sticker database containing 10,648 stickers. The metadata for each sticker includes the original emotion label, the description of emotion, the description of content, and usage recommendations. The sticker database is then evenly distributed into 100 smaller vector databases based on emotional balance.
B.2 Manager Agent
During dataset creation, we limit conversation turns to a maximum of six. The manager agent is programmed to review the conversation every two turns, and to verify the consistency of each sticker used with the conversation context. We observe that with more conversation turns, the likelihood of the agent repeating previous statements increased. Upon detecting such repetition, the manager agent will transition the agent’s LLM to GPT-4 to prevent further repetition.
Model learning rate warmup steps weight decay batch size max length steps
Input PEGS 1e-4 2,000 0.05 128 32 80,000
Output PEGS-Ret 3e-5 4,000 1e-4 36 77 142,220
PEGS-RAG 3e-5 4,000 1e-4 36 77 142,220
PEGS-Gen 1e-5 1,000 0.05 108 32 80,000
Table 6: Hyperparameters for pre-training PEGS.
Model learning rate warmup steps weight decay batch size max length epochs
PEGS-Ret 3e-5 70 0.05 36 77 10
PEGS-RAG 3e-5 70 0.05 36 77 10
PEGS-Gen 5e-5 200 0.05 108 32 4
Table 7: Hyperparameters for fine-tuning the output side of the PEGS on sticker-text pairs.
Model learning rate warmup steps weight decay LoRA-r/alpha/dropout batch size max length epochs
Vicuna-text 5e-4 300 0.00 8/16/0.05 16 4,096 5
Vicuna-tool 5e-4 300 0.00 8/16/0.05 16 4,096 5
ChatGLM3-text 1e-4 300 0.10 8/16/0.05 16 4,096 5
ChatGLM3-tool 1e-4 300 0.10 8/16/0.05 16 4,096 5
PEGS-Ret 3e-6 2,000 0.05 8/16/0.05 4 768 9
PEGS-RAG 3e-6 2,000 0.05 8/16/0.05 4 768 9
PEGS-Gen 1e-4 1,000 0.05 8/16/0.05 4 768 4
Table 8: Hyperparameters for fine-tuning baselines and PEGS on STICKERCONV.
Appendix C More Implementation Details of PEGS
C.1 Pre-training
Multimodal Perception
Our effort is built on BLIP-2 (Li et al., 2023b), which has captured vision-language knowledge from numerous aligned image-text pairs. Zhu et al. (2023) have proven the effectiveness that freezing the pre-trained vision encoder, Q-Former, and LLM, while only pre-training a linear projection layer. We train the input linear projection layer using LAION115M.
Multimodal Generation
Distinguished from realistic images, stickers have an abstract nature and rich emotional expressions. Generic text-to-image models perform well in various scenarios, but they are flawed in sticker generation. Constrained by limited sticker data, there are not enough image-text pairs to turn the base text-to-image model to the style of stickers, and to pre-train for multimodal generation. As a remedy, we adopt a variant of CuteYukiMix [^2] https://civitai.com/models/28169?modelVersionId=163923, a SD model trained on cute cartoon data, as the pre-trained text-to-image model. In the pre-training phase, our objective is to align the representations with the text embeddings, so we train our model on the captions from LAION115M.
The AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate. Table 6 shows the hyperparameter settings in the pre-training phase.
C.2 Fine-tuning
Multimodal Generation
Pre-trained diffusion models (focusing on obvious features, such as characters’ bodies, costumes) are insensitive to sentiment conditions, and it is difficult to directly perform end-to-end fine-tuning to learn emotion elements, we inserted a supervise fine-tuning (SFT) procedure on the image-text pairs with emotion to efficiently cause them to focus on emotion elements. Table 7 shows the hyperparameter settings for this phase.
Joint Learning
We utilize the constructed STICKERCONV for fine-tuning of the entire model to achieve the capabilities of multimodal perception and generation. We train our models in an end-to-end manner. The AdamW optimizer is adopted, and the cosine annealing scheduler is used to adjust the learning rate. Table 8 shows the hyperparameter settings in the joint-learning phase. In this work, $\lambda_{1}=1$, $\lambda_{2}=1$, $\lambda_{3}=1$ in Eq 10.
Appendix D Further Result Analysis
Table 8 presents the hyperparameters for fine-tuning baselines and PEGS. The end-to-end architecture of PEGS necessitates significant graphics memory during training, limiting it to smaller batchsize (4) and maximum lengths (768). Conversely, the baseline model accommodates larger batchsize (16) and maximum lengths (4096). These parameters suggest that PEGS can be further optimized for use on devices with larger graphics memory.
D.1 Analysis of Multimodal Metrics
From Table 2, taking into account the relative frequency of multimodal responses, f-MMr proves to be more reasonable and robust compared to MMr. However, the multimodal response frequency of all models appears to be unusual, with the Vicuna-tool’s frequency being notably low (0.141) and that of other models excessively high ($>$0.8). For comparison, the System’s sticker usage frequency in the STICKERCONV is 0.6. This outcome suggests that future research should focus on adjusting the multimodal response frequency and timing of models to better suit human interaction.
In this paper, we do not use traditional image generation evaluation metrics, such as FID (Heusel et al., 2017), IS (Salimans et al., 2016). These metrics are applied to evaluate the quality and diversity of realistic images. FID and IS are obtained by pre-trained Inception v3 (Szegedy et al., 2016) (pre-trained on realistic images) after certain processing. However, sticker possesses an abstract style and characteristics that do not match the application domains of FID and IS. Based on the above facts, we evaluate the quality and diversity of image generation by human evaluation.
D.2 Analysis of LLM-based Metrics
When using an LLM for scoring, to eliminate randomness, we repeat the scoring for each metric 5 times and calculate the average to determine the final score. From Table 3, the EMP-mm of PEGS outperforms the two tool models, likely owing to PEGS’s end-to-end framework that ensures greater consistency between text and stickers. The top-ranking performance of PEGS-Gen further demonstrates the positive impact of text-sticker relevance on empathy.
D.3 Analysis of Human Metrics
Figure 12 shows the complete questionnaire of human evaluation, which includes the detailed description and scoring criteria for each metric. The five scorers are graduate students in computer science, specializing in research on emotional reasoning or empathetic dialogue, with a thorough understanding of empathy tasks. During scoring, they are unaware of which model serves as the baseline and which is the PEGS.
Metric StiGQ. Et. Es. Con. Flu. Inf. StiD.
ICC 0.920 0.540 0.928 0.457 0.728 0.608 0.365
Table 9: ICC Analysis Results for Human Evaluators
Table 9 shows the results of the inter-rater reliability analysis conducted on the five scorers involved in the manual evaluation, using the Intraclass Correlation Coefficient (ICC) WEIR (2005) as the assessment criterion. Human scorers demonstrate high consistency in metrics like StiGQ, Es, Flu and Inf. The ICC of Et scores are moderate, likely because assessing empathy through text is highly subjective and depends significantly on individual life experiences. Similarly, the moderate ICC scores for Con may stem from the nature of our task as a multimodal empathetic response task. Both the context and response involve multimodal information (sticker), and the text is affected by the stickers. This can, to some extent, disrupt the scorers’ assessment of Con. The ICC of StiD is poor ($<$0.4), primarily due to it being a global metric with a single score per model, leading to higher volatility.
Metrics EMP-txt EMP-mm CON. Rank ID. f-MMr.
Spearman 0.40 0.70 0.87 0.40 0.80 0.90
Kendall 0.40 0.60 0.74 0.40 0.60 0.80
Table 10: Correlation between the metrics and human annotation.
Table 10 illustrates the correlation between various evaluative metrics and human annotations, as evidenced by Spearman’s and Kendall’s correlation coefficients. The EMP-txt metric calculates the correlation with the Et of human metrics, whereas EMP-mm corresponds to the Es. CON is matched with the Con, and Rank correlation is determined by averaging all scores from human evaluations. The ID metric measures correlation with the StiD of human metrics, and f-MMr matches the average of Con, Flu, Inf, and StiD. Both CON and f-MMr exhibit the highest consistencies under two coefficients, achieving 0.87 and 0.90 for Spearman and 0.74 and 0.80 for Kendall, respectively.
These results emphasize the precision and uniformity of these metrics in mirroring human assessments.
EMP-mm and ID metrics also exhibit substantial correlations, suggesting a significant alignment with human perceptions within these domains. Calculating the correlation between the results of LLM-based metrics and the results of human metrics demonstrates that with carefully designed prompts, LLM can effectively evaluate empathy tasks.
In summary, our experiments show that PEGS excels in producing varied and richly informative empathetic text responses, ensuring high contextual consistency, and generating diverse, high-quality stickers to enhance empathetic effects. These findings highlight the benefits of PEGS’s end-to-end architecture. The success of the PEGS demonstrates the effectiveness of Agent4SC and STICKERCONV. Considering the cost of Agent4SC (shown in Appendix G), once PEGS training is completed, inference can be deployed locally to balance cost-effectiveness and efficiency. This paper establishes a valuable resource for future research on multimodal empathetic response generation.
Appendix E Case Study
To further investigate the quality of the multimodal responses predicted by our proposed framework, we show an example of STICKERCONV test data in Figure 7. As we have seen, the textual responses generated by multimodal baselines and our models are consistent with the dialogue context in terms of content and emotion, demonstrating the effectiveness of fine-tuning based on STICKERCONV. All models provide visual responses (stickers) within the given context. The difference is that PEGS has more pronounced emotional tendencies, which facilitates empathy.
To illustrate more straightforwardly the performance of our model in multimodal empathic conversations, we show several cases of user-chatbot interaction. As shown in Figure 8, these cases demonstrate the ability of PEGS to empathize with users and perceive the stickers provided by users, along with generating multimodal responses with vivid stickers. For the positive dialogue context, our model generates positive stickers to enhance interactivity and emotional expression. For the negative dialogue context, our model reassures the user by an appropriate sticker (the hugging behavior in Figure 8 right).
Figure 7: An example of STICKERCONV test set. Top: the multimodal dialogue context between “User” and “System”. Bottom: the multimodal responses generated by our designed multimodal baselines and models.
Figure 8: Examples of conversations by users interacting with PEGS. Users can chat with multimodal content (text and stickers) and will receive multimodal empathetic responses. Left: a conversation characterized by positive emotion (happiness). Right: a conversation characterized by negative emotion (sadness).
Appendix F Prompts
@@ F.1 Sticker Process
F.2 Sticker Agent Chat
Figure 14 shows the chat prompt, the summary encompasses the persona of the user or system, the current status, and the core traits generated based on memory. Figure 15, 17, 18 show the process of used sticker.
F.3 Manager Agent
Figure 16 shows the prompt of manager agent. By using various "format instructions", this prompt can serve multiple purposes, including quality and consistency checks, among others.
F.4 LLM-Based Scorer
@@
Figures 20 depict the templates for the consistency scorer. It is utilized for scoring purely textual responses.
Figures 21 and 24 present the templates for the empathy scorer. The methodology for processing responses of different modalities aligns with the approach described for the consistency scorer.
F.5 Joint Learning
Figure 22 shows the prompt template for joint learning. We employed an instruction template similar to Vicuna v1.5. For every dialogue that uses a sticker, we append an <IMG> identifier.
F.6 Tool Learning
Figure 23 presents the instruction prompts for tool learning. To instruct the LLM in utilizing the SD model for sticker generation, we transform the STICKERCONV to tool format. When the model requires invoking the SD model, it outputs "ASSISTENT Action" along with the corresponding prompt ("ASSISTENT Action Input"). We then feed this prompt into the pre-configured SD model and relay the outcomes back to the LLM as "ASSISTENT Observation".
Appendix G Costing
Extracting knowledge from labeled stickers in Llava using ChatGPT-3.5-turbo, which includes stickers in SER30K dataset, stickers generated by tool models and PEGS, costs approximately \$130.
The construction of 2000 user profiles for Agent4SC with ChatGPT-3.5-turbo incurred a total cost of \$5.5.
Constructing the STICKERCONV dataset with ChatGPT-3.5-turbo incurred a total cost of \$438.53, covering training, validation, and testing sets, at an average of \$0.03 per dialogue session.
According to our test set, evaluating metrics other than rank for each comparison model costs \$22. Combining all model responses with the evaluation Rank metric costs about \$60.
Figure 9: Example Profile of Agent4SC.
Figure 10: Example I from STICKERCONV.
Figure 11: Example II from STICKERCONV.
Figure 12: Human Evaluation Questionnaire.